41.Suppose that )-for o < lowing probabilities: (a) P(l < X) (b) P <X<2.5) (c)P(X= 3) (d) P(X< 4) (e) P3 s x) Determine the fol-
Answer #1

f(x) = e-x for x > 0
(a)
P(X 1)f(r)dr
= 0.368
(b)
P(1 < X < 2.5) = 2.5 | 2.5 f(z)dz
= 0.286
(c)

P(X=3) = 0 ( Point probabilities are zero for continuous distributions)
(d)
P(X<4) = int_{0}^{4}f(x)dx = int_{0}^{4} e^{-x}dx = 1 - rac{1}{e^{4}}
= 0.982
(e)
P(X > 3) | f(x)dx= = ез
= 0.05